# Optimizing Charge Transport Simulation for Deep Learning Enhanced Spatial Resolution of the MÖNCH Detector

### v1

@@ Abstract
Deep learning has the potential to significantly enhance the spatial resolution of hybrid detectors for electrons, as demonstrated for electron microscopy using 200 keV electrons with the MÖNCH 25 µm pitch charge integration pixel detector. The deep learning models have been trained using both simulation data, which are more straightforward to generate with varied parameters, and measurement data, which are more cumbersome to acquire due to the required specialized setup of the electron microscope. However, challenges arise when comparing simulation-based deep learning models to measurement-based models for electrons, as the spatial resolution achieved through simulations is notably inferior to that from measurements. Discrepancies are also observed when directly comparing X-ray simulations with measurements, particularly in the spectral output of single pixels. These observations collectively suggest that current simulations require optimization.
To address this, a Monte Carlo simulation was conducted to model the dynamics of charge carriers within the silicon sensor, aiming to refine charge transport modeling. The simulation encompasses the initial generation of the charge cloud, charge cloud drift, charge diffusion and repulsion, and electronic noise. The simulation results were validated with measurements from the MÖNCH detector for X-rays, and consistencies between the simulation and the measurements were significantly improved by accounting for the charge repulsion.
@@ 1 Introduction
The MÖNCH detector, currently under development at the Paul Scherrer Institut, is a charge-integrating hybrid pixel detector [1]. Its 25 µm pitch, combined with high dynamic range, fast readout, and low noise, demonstrates the potential of hybrid pixel detectors for more electron microscopy applications. Our prior study successfully demonstrated that deep learning can effectively reconstruct the incident position of single electron, thereby enhancing the spatial resolution for 200 keV electrons [2], where performance is primarily constrained by multiple scattering effects. Our deep learning models were trained using both simulation data and measurement data. Simulation data are more straightforward to generate with varied parameters, whereas measurement data acquisition are challenging due to the specialized setup required for the electron microscope. Additionally, simulations offer more detailed information, such as the electron track, which is not available in the measurements.
@@
However, the small pitch of the MÖNCH detector not only enables high spatial resolution but also presents challenges in accurately modeling charge transport. These challenges become apparent when comparing simulation-based deep learning models to those based on measurements. The spatial resolution achieved using simulations is notably inferior to that achieved with measurements. To address the origin of this discrepancy, we conducted comparisons between measurements and simulations generated by [3] using the ProjectionPropagation module for 200 keV electrons and X-rays phohtons.
@@
Figure 1 (a), (b), and (c) show the pixel energy spectra for 200 keV electron clusters and the pixel spectra of 3x3 clusters of copper and silver fluorescence, respectively. At higher energies, the simulation shows a larger number of entries than the measurement, as illustrated in the spectra and amplified in the ratio plots. This discrepancy suggests an underestimation of the charge-sharing effect in the simulation, which requires optimization. Furthermore, when comparing the pixel energy spectra for copper and silver X-rays (Figure 1(b) and (c)), the discrepancies are more pronounced for higher energy depositions.
@@
This study aims to optimize the charge transport simulation for the MÖNCH detector. A time-stepping Monte Carlo simulation is conducted following the charge carrier dynamics in the silicon sensor. The simulation process includes the initial generation of the charge cloud by X-rays or electrons, charge cloud drift, charge diffusion and repulsion, and electronic noise. The charge repulsion, which is currently not considered in established simulation frameworks [3, 4, 5], is proved to be crucial for accurately reproducing the MÖNCH detector measurements. The following sections elaborate on the charge carrier dynamics, simulation algorithm, validation results, discussion, and future outlook.
@@
The dynamics of charge carriers in silicon sensors is governed by the continuity equation:
where represents the charge carrier density, is the diffusion coefficient, denotes the charge carrier mobility, and is the electric field. The generation and recombination of charge carriers are ignored in the depleted sensor volume. Driven by the bias voltage, the charge cloud, consisting of charge carriers, drifts towards the readout electrode. By neglecting charge repulsion, Equation 1.1 simplifies to , with the coordinate system’s origin at the center of the charge cloud. The solution to this equation is a Gaussian distribution with a standard deviation of in one dimension, a charge diffusion formula widely implemented in simulation frameworks.
However, the neglect of charge repulsion in the simulation leads to an underestimation of the charge cloud size, accounting for the discrepancies between the simulation and the measurements in Figure 1. Thus, charge repulsion must be taken into account in the simulation for more accurate modeling of charge transport in the MÖNCH detector. Given that mobility is a complex function of the electric field, finding analytical solutions to the continuity equation that account for charge repulsion is impractical. Therefore, this study developed a time-stepping Monte Carlo simulation to model charge carrier transport in the silicon sensor while incorporating charge repulsion. X-ray photon are simulated in this study due to their point-like energy deposition, which provides an ideal initial condition for simulations. Additionally, the accessibility in measurements using X-ray tubes facilitates direct comparisons.
@@ 2 Time-stepping Monte Carlo simulation
Boundry conditions
The absorption depth of X-rays in the silicon sensor follows an exponential distribution, with the attenuation length determined by the X-ray energy and the sensor material [10]. The absorbed energy is converted into charge at a conversion efficiency of 3.6 eV per electron-hole pair. The number of charge carriers is manually set to 100,000, evenly distributing the total charge. This is because the charge repulsion is the primary focus of this study, and a small number of charge carriers and a sparse charge cloud are sufficient apply the spherical symmetry assumption to calculate the electric field due to the charge repulsion.
The initial charge density distribution adheres to an empirical Gaussian distribution with µm [11], where is the energy deposition. The transport begins at the X-ray absorption depth and ends when the charge cloud center reaches the sensor pixel side.
@@
Holes are collected in the MÖNCH detector. The widely-used Jacoboni-Canali model, which takes into account the saturation of the mobility for high electric fields, is employed to describe the mobility of holes in silicon as a function of the electric field and temperature [6]:
@@
where results from the bias voltage and arises from the charge repulsion, detailed in the following sections. The calculation assumes orthogonality between and , which is valid for charge carriers on the plane across the cloud center and is applied to the entire charge cloud. This assumption maintains spherical symmetry of the charge cloud and simplifies the calculation.
@@
where is the average hole mobility of charge carriers, and is the time step, set to 0.01 ns.
@@
Intermediate results from a simulated copper K X-ray with an absorption position of (0, 0, 2.5 µm) are presented below. Figure 1(a) plots the root mean square (RMS) of charge carrier coordinates over drift time with and without the charge repulsion. The RMS at the end of the simulation including the charge repulsion is 8.46 µm, which is significantly higher than the RMS without the repulsion (7.68 µm).
@@ 3 Validation of the simulation
The validation of the simulation is conducted by comparing the simulation results with measurements obtained using a MÖNCH detector for copper and silver X-ray fluorescence, specifically in terms of the pixel energy spectrum and the charge-weighted center of 3x3-pixel clusters. X-ray clusters of both the measurements and the simulation adher to the same selection criteria and processing steps.
Forming 3x3-pixel clusters
@@
The simulation previously described is repeated for various absorption depths every 5 µm. One million X-rays are simulated, with the absorption depth sampled from an exponential distribution corresponding to the respective attenuation length and a random incident position uniformly distributed in the plane. For each X-ray, the charge carrier distribution relative to the cloud center is sampled from the fitted generalized Gaussian distribution for the nearest simulated absorption depth. Pixel noise, approximately as determined from measurements, is added to each pixel. The 3x3-pixel clusters are centered on the pixel with the maximum charge. Additionally, K X-rays are simulated following the same procedure, with the intensity ratio to K X-rays derived from an empirical study [7] for copper. In the case of the silver K the ratio was obtained from the double-Gaussian fit to K and K measurement.
@@
Measurements were conducted using the MÖNCH detector with an X-ray tube featuring copper and silver targets. To process the raw data from the charge-integration MÖNCH detector, dark current subtraction was applied. Pixel-wise gain calibration was performed to convert the pixel readout from Arbitrary Digital Units to energy. The 3x3-pixel clusters were centered on the pixel with the maximum energy. Pile-up clusters were excluded from the analysis by ensuring no neighboring pixels of the cluster belonged to another cluster.
@@
An 1 keV cluster energy selection around the K peak was applied to both data sets, further excluding noise and pile-up clusters.
Pixel energy spectrum
@@
The pixel energy spectra of the MÖNCH detector for selected copper and silver X-ray clusters are illustrated in Figure 3 (a) and (b), respectively. Both figures show good agreement between the simulation and the measurements, with some remaining discrepancies. The simulation results exhibit a higher number of entries in lower energy bins, while the measurements show a higher number of entries in higher energy bins especially at the K peak. These discrepancies are primarily attributed to noise modeling, the non-linearity of the charge amplification, and cross-talk between the pixels, which causes a bump in the K peak visible only in the measurements. Overall, significant improvements are observed in the simulation results compared to results obtained without accounting for charge repulsion by comparing Figure 3 and Figure 1.
Charge weighted center ,
@@
The charge weighted centers and are widely used in position interpolation to improve spatial resolution [8, 9]. For the 3x3-pixel clusters, the charge weighted centers and are calculated as follows:
@@
Here denotes the energy of the pixel111Charge and energy are translated using a conversion coefficient of 3.6 eV per electron-hole pair. at the -th row and the -th column. With the origin of the coordinate system defined as the lower left corner of the center pixel, the charge weighted centers and range from 0 to 1.
@@
The 2D histograms of the charge weighted centers and for copper and silver X-rays are depicted in Figures 4(a) and (b,) respectively. The z-axis in logarithmic scale represents the number of entries. The spectra after projection, obtained from both measurements and simulations, are displayed in Figure 4(c) and (d). The simulation aligns well with the measurements for copper X-rays, while there remain some discrepancies for silver X-rays. The sources of these discrepancies will be detailed in the following section.
@@ 4 Discussion, conclusion and outlook
The 25 µm pitch of the MÖNCH detector presents the challenge of a more accurate modeling of charge transport, while simultaneously offering a unique platform for the investigation. Meanwhile, we should note that the difference in the charge distribution with and without repulsion, as exampled in Figure 1(a), is less noticeable for hybrid pixel detectors with a larger pixel pitch, such as 55 µm of the Medipix detectors [12] and 75 µm of the Jungfrau detectors [13].
@@
While incorporating the charge repulsion into the simulation has significantly improved the agreement between the simulation and the measurements, some discrepancies persist. The simulation assumes a spherically symmetric charge distribution to simplify the calculation of the electric field due to charge repulsion, an assumption which does not strictly hold. Additionally, the gradient of the drift electric field is neglected in the simulation. When the charge cloud approaches the readout electrode, the symmetry of the charge cloud is broken. Furthermore, up to a 10% deviation in the mobility model exists [6].
@@
Further research is necessary to implement this approach for electron simulations. Modeling the charge cloud for electrons is more challenging than for X-rays due to the continuous energy deposition spectrum. The parameters of the charge cloud distribution should be modeled as a function of both the deposition depth and the deposition energy. Additionally, the repulsion between charge clouds must be taken into account. With optimized simulations, higher-quality simulation samples can be generated for deep learning applications. Various design parameters of the MÖNCH detector can be investigated to exploit the charge sharing effect.

### v3

# Optimizing Charge Transport Simulation for Hybrid Pixel Detectors

@@ Abstract
To enhance the spatial resolution of the MÖNCH 25 µm pitch hybrid pixel detector, deep learning models have been trained using both simulation and measurement data. Challenges arise when comparing simulation-based deep learning models to measurement-based models for electrons, as the spatial resolution achieved through simulations is notably inferior to that from measurements. Discrepancies are also observed when directly comparing X-ray simulations with measurements, particularly in the spectral output of single pixels. These observations collectively suggest that current simulations require optimization.
To address this, the dynamics of charge carriers within the silicon sensor have been studied using Monte Carlo simulations, aiming to refine the charge transport modeling. The simulation encompasses the initial generation of the charge cloud, charge cloud drift, charge diffusion and repulsion, and electronic noise. The simulation results were validated with measurements from the MÖNCH detector for X-rays, and the agreement between measurements and simulations was significantly improved by accounting for the charge repulsion.
@@ 1 Introduction
The MÖNCH detector, currently under development at the Paul Scherrer Institute, is a charge-integrating hybrid pixel detector [1]. Its 25 µm pitch, combined with high dynamic range, fast readout, and low noise, makes the MÖNCH detector a promising candidate for electron microscopy applications. Our prior study [2] demonstrated that deep learning can effectively reconstruct the incident position of single electrons, thereby enhancing the spatial resolution for 200 keV electrons, in the case of which performance is primarily constrained by multiple scattering effects. The deep learning models were trained using both simulation data and measurement data. Simulation data are more straightforward to generate with varied parameters, whereas measurement data acquisition is challenging due to the specialized setup required for the electron microscope. A primary challenge with measurement data lies in accurately determining the impact point, compounded by issues such as beam current stability and precise calibration. Additionally, simulations offer more detailed information, such as the electron track, which is not available in the measurements.
@@
However, the small pitch of the MÖNCH detector not only enables high spatial resolution but also presents challenges in accurately modeling charge transport. These challenges become apparent when comparing simulation-based deep learning models to those trained on measurements. In particular, The spatial resolution achieved using simulations is notably inferior to that achieved with measurements. To address the origin of this discrepancy, we compared measurements and simulations generated by Allpix Squared [3] for 200 keV electrons and X-rays photons. Simulations were conducted using the ProjectionPropagation module with employing the MÖNCH detector geometry and measurement conditions: a sensor thickness of 320 µm with a 25 µm pitch, a depletion of 30 V and a bias voltage of 90 V for collecting holes, respectively, and a temperature of 293 K.
@@
Figure 1 (a), (b), and (c) show the single pixel energy spectra for 200 keV electrons and copper and silver fluorescence X-rays ( keV and keV), respectively. The selected pixels are those with energy over for electrons and from 33-pixel clusters for X-ray photons. At higher energies, the simulation shows a larger number of entries than the measurement, as illustrated in the spectra and amplified in the ratio plots. This discrepancy suggests an underestimation of the charge-sharing effect in the simulation, which requires optimization. Furthermore, when comparing the pixel energy spectra for copper and silver X-ray fluorescence (Figure 1(b) and (c)), the discrepancies are more pronounced for higher energy depositions.
@@
This study aims to optimize the charge transport simulation for the highly pixelated MÖNCH detector. A time-stepping Monte Carlo simulation is conducted following the charge carrier dynamics. The simulation process includes the initial generation of the charge cloud by X-rays or electrons, charge cloud drift, charge diffusion and repulsion, and electronic noise. The charge repulsion, which is currently not considered in established simulation frameworks [3, 4, 5], is proved to be crucial for accurately reproducing the MÖNCH detector measurements. The following sections elaborate on the charge carrier dynamics, simulation algorithm, validation results, discussion, and future outlook.
@@
The dynamics of charge carriers in silicon sensors are governed by the continuity equation:
where represents the charge carrier density, is the diffusion coefficient, denotes the charge carrier mobility, is the Boltzmann constant, is the temperature, denotes the unit charge, and is the electric field. The generation and recombination of charge carriers are ignored in the depleted sensor volume. Driven by the bias voltage, the charge cloud, consisting of charge carriers, drifts towards the readout electrode. By neglecting charge repulsion, Equation 1.1 simplifies to , with the coordinate system’s origin at the charge cloud center at . The solution to this equation is a Gaussian distribution with a standard deviation of in one dimension, a charge diffusion formula widely implemented in simulation frameworks.
However, neglecting charge repulsion in the simulation leads to an underestimation of the charge cloud size, accounting for the discrepancies with measurements in Figure 1. To improve the accuracy in modeling charge transport in the MÖNCH detector, charge repulsion must be included in the simulation. Given that mobility is a complex function of the electric field, it is impractical to find analytical solutions to the continuity equation that account for charge repulsion. Therefore, this study implemented a time-stepping Monte Carlo simulation to model charge carrier transport in the silicon sensor while incorporating charge repulsion. X-ray photons are studied due to their point-like energy deposition, which provides an ideal initial condition for simulations. Additionally, the ease of in-air measurements with metal X-ray fluorescence facilitates a direct comparison between simulations and measurements.
@@ 2 Time-stepping Monte Carlo simulation
Boundary conditions
The absorption depth of X-rays in the silicon sensor follows an exponential distribution, with the attenuation length determined by the X-ray energy and the sensor material [10]. The absorbed energy is converted into charge at an efficiency of 3.6 eV per electron-hole pair.
The initial photonelectron follows an semiempirical distribution with a Bethe range [11, 12] given by µm, where g/cm3 is the silicon density and is the energy deposition in keV. This expression is valid within the energy deposition range of 5 to 25 keV. The 1D standard deviation of the corresponding Gaussian profile is given by µm [13]. The transport begins at the X-ray absorption depth and ends when the charge cloud center reaches the sensor surface driven by the bias voltage.
@@
The MÖNCH detector operates in hole-collection mode. The widely-used Jacoboni-Canali model, which takes into account the saturation of the mobility for high electric fields, is employed to describe the mobility of holes in silicon as a function of the electric field and temperature [6]:
@@
where results from the bias voltage and arises from the charge repulsion, detailed in the following sections. The calculation assumes orthogonality between and , which is valid for charge carriers on the - plane across the cloud center and is applied to the entire charge cloud. This assumption maintains spherical symmetry of the charge cloud and simplifies the calculation.
@@
where is the average hole mobility of charge carriers and is the time step, set to 0.01 ns.
@@
Intermediate simulation results from 8.05 keV X-rays (i.e., Cu K X-fluorescence energy) with an absorption position of (0, 0, 2.5 µm) are presented below. The Monte Carlo simulation is repeated multiple times to achieve a statistical power of approximately 100,000 charge carriers. Figure 1(a) plots the root mean square (RMS) of charge carrier coordinates over drift time with and without the charge repulsion. The 1D RMS width at the end of the simulation, including the charge repulsion, is 8.46 µm, which is significantly higher than the 1D RMS without the repulsion (7.70 µm).
@@ 3 Validation of the simulation
The validation of the simulation is conducted by comparing the simulation results with measurements obtained using a MÖNCH detector for copper and silver X-ray fluorescence, specifically in terms of the pixel energy spectrum and the charge-weighted center of 33-pixel clusters. X-ray clusters of both the measurements and the simulation follow the same selection criteria and processing steps.
Forming 33-pixel clusters
@@
The previously described simulation is repeated for various absorption depths, with steps of 1 µm. Then, one million X-rays are simulated, with the absorption depth sampled from an exponential distribution corresponding to the respective attenuation length and a random incident position uniformly distributed in the - plane. For each X-ray photon, the charge carrier distribution relative to the cloud center is sampled from the fitted generalized Gaussian distribution for the nearest simulated absorption depth. Pixel noise, including electronic noise (approximately as determined from measurements) and Poisson fluctuations of the charge carriers, is added to each pixel. The 33-pixel clusters are centered on the pixel with the maximum charge. Additionally, K X-rays are simulated following the same procedure, with the intensity ratios to K X-rays given by [7].
@@
Measurements were conducted using the MÖNCH detector with an X-ray tube featuring copper and silver targets. To process the raw data from the charge-integration MÖNCH detector, dark current subtraction was applied. Pixel-wise gain calibration was performed to convert the pixel readout from Arbitrary Digital Units to energy. The 33-pixel clusters were centered on the pixel with the maximum energy. Pile-up clusters were excluded from the analysis by ensuring that no neighboring pixels of the cluster belonged to another cluster.
@@
A cluster energy selection of keV of the K peak was applied to both data sets to further exclude noise and pile-up clusters.
Simulated and measured pixel energy spectra
@@
The pixel energy spectra of the MÖNCH detector for selected copper and silver X-ray clusters are illustrated in Figure 3 (a) and (b), respectively. Both figures demonstrate good agreement between the simulation and the measurements, though some discrepancies remain. The simulation results show a higher number of entries in the first half of the energy range, while the measurements exhibit more entries in the second half. These differences are primarily attributed to the simplified noise modeling, as well as non-linearity in charge amplification and pixel cross-talk, which are not accounted for in the simulation. Overall, significant improvements are observed in the simulation results comparing results obtained without (Figure 1), and with (Figure 3) charge repulsion.
Simulated and measured charge weighted center , parameter distributions
@@
The charge weighted centers and are widely used in position interpolation to improve spatial resolution [8, 9]. For the 33-pixel clusters, the charge weighted centers and are calculated as follows:
@@
Here denotes the energy measured by the pixel111Charge and energy are translated using a conversion coefficient of 3.6 eV per electron-hole pair. at the -th row and the -th column. With the origin of the coordinate system defined as the lower left corner of the center pixel, the charge weighted centers and range from 0 to 1.
@@
The 2D histograms of the charge weighted centers and for copper and silver X-rays are depicted in Figures 4(a) and (b,) respectively. The color bar in logarithmic scale represents the number of entries. The spectra after projection, obtained from both measurements and simulations, are displayed in Figure 4(c) and (d). The simulation aligns well with the measurements for copper X-rays, while there remain some discrepancies for silver X-rays. The possible sources of these discrepancies are detailed in the following section.
@@ 4 Discussion, conclusion and outlook
The 25 µm pitch of the MÖNCH detector presents the challenge of a more accurate modeling of charge transport, while simultaneously offering a unique platform for the investigation of this phenomenon. Meanwhile, we should note that the difference in the charge distribution with and without repulsion, as exampled in Figure 1(a), is less noticeable for hybrid pixel detectors with a larger pixel pitch, such as 55 µm of the Medipix detectors [14] and 75 µm of the Jungfrau detectors [15].
@@
While incorporating the charge repulsion into the simulation has significantly improved the agreement between the simulation and the measurements, some discrepancies persist. The simulation assumes a spherically symmetric charge distribution to simplify the calculation of the electric field due to charge repulsion, an assumption which does not strictly hold. Additionally, the gradient of the drift electric field within the charge cloud is neglected in the simulation. When the charge cloud approaches the readout electrode, the symmetry of the charge cloud is broken. Furthermore, mobility uncertainty is up to 10% in the Jacoboni-Canali model [6], which may contribute to the discrepancies.
When comparing the results of the copper and silver X-rays, the discrepancies are more pronounced for the silver X-rays. Note that the impact of charge repulsion mainly occurs within the first few nanoseconds and then saturates (see Figure 2). The saturation time also influences the charge distribution. For silver X-rays, more photons are absorbed in the deeper sensor volume, which amplifies the effect mentioned above. One possible reason for the differing saturation times is the absence of electrons, the other type of charge carrier, in the simulation, which accelerates the development of the repulsion effect.
@@
Further research is necessary to implement this approach for electron simulations. Modeling the charge cloud for electrons is more challenging than for X-rays due to the continuous energy deposition spectrum. The parameters of the charge cloud distribution should be modeled as a function of both the deposition depth and the deposition energy. Additionally, the repulsion between charge clouds must be considered. With optimized simulations, higher-quality simulation samples can be generated for deep learning applications. Various design parameters of the MÖNCH detector, e.g., the sensor thickness and the bias voltage, can be investigated to exploit the charge sharing effect for a better spatial resolution.
Code availability
TThe simulation source code, along with an example Jupyter notebook that includes fitting and plotting scripts, as well as histograms of the measurement results, can be accessed at https://github.com/slsdetectorgroup/ChargeTransportSimulation.
Acknowledgements
The authors would like to express their gratitude to Dr. Emiliya Poghosyan and Dr. Elisabeth Müller for their expertise in conducting measurements with the electron microscope.